### v1

# Random sampling of partitions and contingency tables: Two practical examples of the Burnside process

Abstract
The Burnside process is a general algorithm for sampling a uniformly chosen orbit of a finite group $G$ acting on a finite set $\mathscr{X}$. For example, if $\mathscr{X}=G$ and $G$ acts on itself by conjugation ($s^{t}=t^{-1}st$), then the orbits are conjugacy classes. When $G$ is the symmetric group $S_{n}$, the conjugacy classes are indexed by partitions of $n$, so the Burnside process gives a way to sample partitions. If $\mathscr{X}=S_{n}$ and $G$ is a product of symmetric groups, then the orbits are labeled by contingency tables: non-negative integer arrays with given row and column sums. Actually carrying out the Burnside process requires new combinatorics and group theory. This is worked out and illustrated for these two examples. For partitions, we also developed a new Markov chain called the reflected Burnside process which greatly improves the mixing of the Burnside process.
1 Introduction
Let ${\mathscr{X}}$ be a finite set and $G$ a finite group acting on ${\mathscr{X}}$. This splits ${\mathscr{X}}$ into disjoint orbits
${\mathscr{X}}=\mathscr{O}_{1}\sqcup\mathscr{O}_{2}\sqcup\cdots\sqcup\mathscr{O}_{Z}.$
For $x\in{\mathscr{X}}$, $\mathscr{O}_{x}$ denotes the orbit containing $x$. Natural questions about the orbits are
  • How many orbits are there?
  • How large is a typical orbit?
  • How can an orbit be uniformly chosen?
Solving the third problem helps with the first two. The Burnside process [39] is a Markov chain on ${\mathscr{X}}$. A single step of this Markov chain transitions from $x$ to $y$ as follows:
  1. 1.
    From $x\in{\mathscr{X}}$ uniformly choose $g\in G_{x}=\{g:x^{g}=x\}$.
  2. 2.
    From $g\in G$, uniformly choose $y\in{\mathscr{X}}_{g}=\{y:y^{g}=y\}$.
The probability of moving from $x$ to $y$ is therefore
$P(x,y)=\frac{1}{|G_{x}|}\sum_{g\in G_{x}\cap G_{y}}\frac{1}{|{\mathscr{X}}_{g}|}.$ (1)
This $P$ is an ergodic reversible Markov kernel with stationary distribution
$\pi(x)=\frac{Z^{-1}}{|\mathscr{O}_{x}|},\quad Z=\#\text{ of orbits}.$ (2)
Thus, ‘running the Burnside process’ and reporting the orbit each time gives a way of choosing a uniform orbit.
Example 1.1.
Let ${\mathscr{X}}=C_{2}^{n}$ be the set of binary $n$-tubles. Let $G=S_{n}$ act on ${\mathscr{X}}$ by permuting coordinates. There are $n+1$ orbits with
$\mathscr{O}_{i}=\{x:|x|=i\}\text{ for }0\leq i\leq n,$
where $|x|$ is the number of $1$’s in $x$. The Burnside process proceeds as follows:
  1. 1.
    From $x$, choose $\sigma\in S_{n}$ fixing $x$. Such a permutation must permute the $1$’s and 00’s amongst themselves. Thus, if $|x|=i$, then $\sigma\in G_{x}\cong S_{i}\times S_{n-i}$. It is easy to sample from $S_{i}\times S_{n-i}$ and hence $G_{x}$.
  2. 2.
    From $\sigma$, we must choose $y\in C_{2}^{n}$ fixed by $\sigma$. This is also easy to do. First, decompose $\sigma$ into disjoint cycles $\sigma=c_{1}c_{2}\cdots c_{j}$. Then, label each cycle 00 or $1$ with probability $1/2$ independently. The label of $y_{i}$ is determined by the cycle containing $i$.
The Burnside process is reviewed in more detail in Section 2. For more general ${\mathscr{X}}$ and $G$, it can be challenging to carry out the two steps above. In Section 3, we study the Burnside process for integer partitions of $n$. Here, $S_{n}$ acts on itself by conjugation. Empirically, we find that adding ‘deterministic jumps’ to the Burnside process creates a Markov chain that mixes rapidly. In Section 4, an implementable method is developed for general double coset spaces $H\backslash G/K$. This is specialized to contingency tables in Section 5. Both Section 5 and 3 are self-contained with literature reviews, algorithms and worked examples. They may be read now for further motivation. The final section summarizes practical advice. Our key finding is that the ‘lumped chain’ chain can be enormously more efficient for large problems. It also points to new applications where the present approaches should be applicable. Our implementation of the Burnside process for partitions and contingency tables is available in Julia [10] at https://github.com/Michael-Howes/BurnsideProcess/. The repository also contains notebooks for recreating the figures in this article.
2 Background on the Burnside process
The Burnside process was introduced by [39] and developed for ‘computational Pólya theory’ in [36, 35]. They show that, sometimes, sampling is ‘hard’ and the Burnside process can be exponentially slow (for special examples). The present paper shows, for some examples, the Burnside process can be extremely efficient. See [39, 1] for connections between the Burnside process and the rapidly mixing Swendsen–Wang algorithm.
In [22] it is shown that the Burnside process can be ‘lumped to orbits’. That is, run directly on the orbit space. This orbit chain is symmetric and so has a uniform stationary distribution. Discerning useful formulae or descriptions for the chance of transitioning from one orbit to another can take work (!).
Example 2.1 (Example 1.1 continued).
For ${\mathscr{X}}=C_{2}^{n}$ and $G=S_{n}$, the orbits are indexed by $\{0,1,\ldots,n\}$. Let $\overline{P}(j,k)$ be the change of moving from $j$ to $k$ in one step of the lumped Burnside process. The following result was used in [22] to bound the mixing time of the Burnside process for this example.
Proposition 2.2 (Section 3 in [22]).
Let
$\alpha^{n}_{k}=\binom{2k}{k}\binom{2(n-k)}{n-k}/2^{2n}\quad 0\leq k\leq n,$
be the discrete arcsine distribution [31, Chapter 3]. Then
$\overline{P}(0,k)$ $=\overline{P}(n,k)=\alpha^{n}_{k},$
$\overline{P}(j,k)$ $=\sum_{l}\alpha^{j}_{l}\alpha^{n-j}_{k-l}\quad(j-k-n)_{+}\leq l\leq\min(j,k),$ (3)
$\overline{P}(j,k)$ $=\overline{P}(k,j)=\overline{P}(n-j,k)=\overline{P}(j,n-k)\text{ for all }k,j.$
Proposition 2.2 gives a new way to run the lumped Burnside chain. Equation 2.2 says that the distribution $\overline{P}(j,\cdot)$ is the convolution of two discrete arcsine distributions. Thus, to transition from $j$ to $k$ we can independently sample $k_{1}$ and $k_{2}$ from the discrete arcsine distributions with parameters $j$ and $n-j$ and then set $k=k_{1}+k_{2}$. It is easy to sample from the discrete arcsine distribution since the discrete arcsine distribution is a Beta-Binomial distribution with parameters $n$ and $\alpha=\beta=1/2$.
The Burnside process has been usefully implemented for a number of special cases.
Example 2.3 (Random matrices).
Let $U_{n}(q)$ be the group of uni-upper triangular matrices with entries in a finite field over $\mathbb{F}_{q}$. When $n$ is at all large (e.g. $n=20$) the conjugacy classes of $U_{n}(q)$ are unknown and, in a sense, ‘unknowable’. See [24] for details. The Burnside process gives a way of sampling a uniform conjugacy class. The equivalence between sampling and counting [41, 13] gives a route to effective enumeration of $|C_{n}(q)|$–the number of conjugacy classes. This program has be carried out in [30] who get useful estimate of $|C_{n}(2)|$ for $n$ up to $40$.
Here ${\mathscr{X}}=G=U_{n}(q)$. The Markov chain runs on ${\mathscr{X}}$. For the first step, from a matrix $M\in U_{n}(q)$ one must sample $M^{\prime}\in U_{n}(q)$ so that $MM^{\prime}=M^{\prime}M$. For fixed $M$ this is a linear algebra problem over $\mathbb{F}_{q}$. Gaussian elimination can be used to find a basis for the solution space. Choosing a random linear combination of basis vectors gives a uniform choice of $M^{\prime}$. The second step is the same as the first (choose $M^{\prime\prime}$ with $M^{\prime}M^{\prime\prime}=M^{\prime\prime}M^{\prime}$). Of course, this is a serious undertaking, but it is quite possible and has been usefully carried out [30]. It provides an important supplement (and check) on exact computation.
Example 2.4.
Let $\mathscr{C}_{n}$ be the set of labeled trees on $n$ vertices rooted at the vertex 1. Cayley’s formula shows $|\mathscr{C}_{n}|=n^{n-2}$ (so $|\mathscr{C}_{4}|=16$, see Figure 1). The symmetric group $S_{n-1}$ acts on $\mathscr{C}_{n}$ by permuting the vertices and fixing the roots. The orbits of $S_{n-1}$ on $\mathscr{C}_{n}$ are unlabeled Pólya trees $\mathscr{T}_{n}$ (see $\mathscr{T}_{4}$ in Figure 2). There is no formula for $|\mathscr{T}_{n}|$ and enumerative questions are challenging [25].
Consider the Burnside process with ${\mathscr{X}}=\mathscr{C}_{n}$ and $G=S_{n-1}$. The two steps are:
  1. 1.
    From $t\in\mathscr{C}_{n}$, choose $\sigma\in S_{n-1}$ uniformly with $t^{\sigma}=t$.
  2. 2.
    From $\sigma$, choose $t_{1}$ uniformly with $t_{1}^{\sigma}=t_{1}$.
Both steps are challenging, in theory and practice. [9] successfully implement the Burnside process. This requires an extension of Cayley’s formula to a formula for the number of labeled trees fixed by a given permutation, and uses substantial input from the computational graph theory community [6, 7] (and their accompanying software https://automorphisms.org/). The implementation efficiently yields samples from trees of size 10,000.
1234124313241342142314321234132414231234124313241342142314321234
Figure 1: The $16$ labeled rooted trees at $1$, for $n=4$. Here and below the root vertex is indicated as solid, and non-root vertices are hollow. Reproduced from [9].
Figure 2: The $4$ unlabeled rooted trees, for $n=4$. Reproduced from [9].
It is natural to ask about rates of convergence for the Burnside process. In Section 3.3 we find empirically that, when combined with an involution, the Burnside process on partitions mixes incredibly rapidly. For Pólya trees as described in Example 2.4, [9] report empirical convergence in 30 steps for $n=10,000$ (!).
There has been some work on rates of convergence in special cases.
Example 2.5 (Example 1.1 revisited).
For ${\mathscr{X}}=C_{2}^{n}$ and $G=S_{n}$, [29] show that
Theorem 2.6 (Theorem 1.1 in [29]).
Let $x_{0}=(0,0,\ldots,0)$, and $P_{x_{0}}^{j}$ be the probability distribution on ${\mathscr{X}}$ induced by starting at $x_{0}$ and taking $j$ steps of the Burnside process. Then, for all $j\geq 4$ and all $n$
$\frac{1}{4}\left(\frac{1}{4}\right)^{j}\leq\|P_{x_{0}}^{j}-\pi\|_{\mathrm{TV}}\leq 4\left(\frac{1}{4}\right)^{j}.$
This shows that a finite number of steps suffices for convergence no matter how large $n$ is. The proof uses an explicit diagonizational of the lumped chain $\overline{P}$ where all the eigenvalues and eigenvectors are known. For the analogous chain on $C_{k}^{n}$, Aldous [3, Section 12.1.6] gives a coupling argument that shows that order $k\log n$ steps are sufficient. Zhong [47] studies a weighted Burnside process on $C_{k}^{n}$ and provides a sharper analysis.
The Burnside process on contingency tables is studied in Chapter 5 of [23] where he shows that the Burnside process mixes quickly for large, sparse tables. We also study this example in Section 5. Our emphasis is on implementing the lumped Burnside process and on the connection between contingency tables and double cosets. Our results thus nicely complement [23].
Careful analysis of a Burnside process for set partitions in is [44]. Mixing time results for a nice collection of conjugacy class chains is in [45]. The results in [45] work for ‘CA groups’. A group is a CA group if the centralizer of any non-central element is Abelian. These include the Heisenberg group mod $p$, the affine group mod $p$ and $\mathrm{GL}_{2}(\mathbb{F}_{p})$. For these groups, a combination of character theory, coupling and geometric arguments give matching upper and lower bounds.
All of this said, careful mixing time analysis of the Burnside process is in an embryonic state. This includes the two main examples below. We are working on it.
3 Partitions
The commuting graph process on a finite group $G$ is an instance of the Burnside process. Here $\mathscr{X}=G$ and $G$ acts on itself by conjugation $x^{g}=g^{-1}xg$. The orbits are conjugacy classes of $G$ and the Burnside process gives a way of sampling a uniform conjugacy class. Here both steps of the Burnside process are the same. From $x$, first sample $g$ that commutes with $x$ and then sample $y$ that commutes with $g$. The commuting graph process on $G$ is just a single step: from $x$, sample $y$ that commutes with $x$.
In the symmetric group $S_{n}$, the conjugacy class of a permutation is determined by its cycle type. This means that the conjugacy class in $S_{n}$ are in bijection with $\mathscr{P}_{n}$–the set of integer partitions of $n$. The commuting graph process on $S_{n}$ therefore gives a method for uniformly sampling integer partitions. Like all Burnside processes, the commuting graph walk can be lumped to a Markov chain on $\mathscr{P}_{n}$. We show that the lumped process can be implemented efficiently and has a per-step complexity on the order of $\sqrt{n}(\log n)^{2}$. We also implement a second Markov chain that transposes the partition in-between steps of the lumped Burnside process. This ‘reflected Burnside process’ empirically converges very quickly and has a similar per-step complexity. Empirically, less than 20 steps are need for convergence for $n$ as large as $10^{8}$.
Section 3.1 gives some background on uniform partitions. It surveys existing methods for sampling uniform partitions, and some known results about what large uniform partitions “look like”. Section 3.2 describes the lumped Burnside process for partitions and contains a derivation of the average per-step complexity. The reflected Burnside process is described in Section 3.3 which also contains the results of our empirical analysis of the mixing time. The reflected Burnside process is an example of speeding up a Markov chain by adding ‘deterministic jumps’. Section 3.3 surveys the literature on this technique. We are excited to have found an example ‘in the wild.’
3.1 Uniform partitions
As before let $\mathscr{P}_{n}$ denote the set of integer partitions of $n$. We will represent partitions in exponential notation as $1^{a_{1}}2^{a_{2}}\cdots n^{a_{n}}$ or $(a_{l})_{l=1}^{n}$ where $a_{l}$ is the number of parts of size $l$. This notation is useful for describing the lumped Burnside process in Section 3.2. The set of integer partitions of $n$ is thus
$\mathscr{P}_{n}=\bigg{\{}(a_{l})_{l=1}^{n}:a_{l}\in\mathbb{Z}_{\geq 0},\sum_{l\geq 1}la_{l}=n\bigg{\}}.$
Chapter 10 of [43] contains an algorithm for uniformly sampling for $\mathscr{P}_{n}$. This algorithm is based on a bijective proof of an identity of Euler’s that relates partitions of $n$ to partitions of $m<n$ and divisors of $m-n$. Their algorithm requires computing and storing $p(m)=|\mathscr{P}_{m}|$ for all $m\leq n$ which can be done efficiently by the Pentagonal Number Theorem – another identity of Euler.
There are also a number of algorithms for sampling partitions that are based on rejection sampling [34, 2]. These methods generate a partition of a random integer $N$ such that conditional on $N=n$, the partition is uniformly distributed on $\mathscr{P}_{n}$. To get a partition of size $n$, these methods repeated sample partitions until one is observed with $N=n$. These methods fit into a broader class of algorithms for sampling from combinatorial structures called Boltzmann Samplers [20, 32]. The efficiency of the rejection sampling methods thus depends on the probability of $N=n$. For the method presented in [34], this probability is on the order of $n^{-3/4}$. While for the probabilistic divide and conquer method in [2], the probability that $n=N$ is bounded away from 00 as $n\to\infty$. The complexity of the method in [2] is $O(n^{1/2+\varepsilon})$ for any $\varepsilon>0$.
[34] contains many results about the asymptotic distribution of ‘features’ of random partitions as $n\to\infty$. In Section 3.3, these asymptotic results are used to empirically assess convergence of the reflected Burnside process. We will use the following theorems from [34].
Theorem 3.1 (Theorem 2.2 in [34]).
Let $(a_{l})_{l\geq 1}$ be a uniformly distributed partition of $n$. Then for fixed $l$ and all $x\geq 0$
$\lim_{n\to\infty}\mathbb{P}\left(\frac{\pi}{\sqrt{6n}}lX_{l}\leq x\right)=1-e^{-x}.$
Theorem 3.2 (Theorem 2.3 in [34]).
Let $(a_{l})_{l\geq 1}$ be a uniformly distributed partition of $n$. Let $I=\sum_{l\geq 1}a_{l}$ be the number of parts in $(a_{l})_{l\geq 1}$, then for all $x\in\mathbb{R}$
$\lim_{n\to\infty}\mathbb{P}\left(\frac{\pi}{\sqrt{6n}}I-\log\frac{\sqrt{6n}}{\pi}\leq x\right)=e^{-e^{-x}}.$
Stated probabilistically, Theorem 3.1 states that $a_{l}$ converges to an exponential random variable when appropriately normalized. Likewise, Theorem 3.2 states that $I$ converges to a Gumble distribution when appropriately normalized. Theorem 3.1 is actually a weaker statement than Theorem 2.2 in [34] which gives the limiting distribution of several entries $(a_{1},\ldots,a_{l_{n}})$ for $l_{n}$ growing at a rate slower than $n^{1/4}$. Also, Theorem 2.3 in [34] is stated in terms of the largest part of $(a_{l})_{l\geq 1}$. However, as noted in the introduction of [34], the largest part has the same distribution as the number of parts. This can be seen be reflecting the Ferrers diagram of the partition (see Figure 1.1 in [34]). Figure 3 contains an illustration of Theorems 3.1 and 3.2, and our new algorithm. The limit theory is in accord with finite $n$ “reality”.
(a) Number of ones ($n=10^{4}$)
(b) Number of ones ($n=10^{6}$)
(c) Number of parts ($n=10^{4}$)
(d) Number of parts ($n=10^{6}$)
Figure 3: Comparisons between samples from the reflected Burnside process and the asymptotic distributions in Theorems 3.1 and 3.2. The histogram represents $10^{5}$ partitions generated from the reflected Burnside process (Section 3.3). Each partition was generated by taking $20$ steps of the reflected Burnside process initialized at the partition $1^{n}$. The dark lines show the asymptotic densities of the two features. Figures 3(a) and 3(b) show that the number of ones is well approximated by an exponential distribution for both $n=10^{4}$ and $n=10^{6}$. Figure 3(c) shows some disagreement between the number of parts and the limiting Gumble distribution and the fit is much better for $n=10^{6}$. Figures 3(b) and 3(d) also suggest that the reflected Burnside process reaches its stationary distribution in just 20 steps for large $n$. On a personal computer, it took 5 minutes to sample all $10^{5}$ partition when $n=10^{4}$. For $n=10^{6}$, the same computer took just under an hour.
3.2 The Burnside process for partitions
This section describes our implementation of the lumped Burnside process on $\mathscr{P}_{n}$. We will see that the lumping reduces the complexity of the Burnside process from order $n$ to order $\sqrt{n}(\log n)^{2}$.
In this setting, the unlumped process is the commuting graph walk on $S_{n}$. The lumped process is the induced Markov chain on $\mathscr{P}_{n}$ given by mapping a permutation to its cycle type. Thus, to run the lumped Burnside process we need a description of the cycle type of $\tau$ when $\tau$ is uniformly sampled from $C_{S_{n}}(\sigma)$–the centralizer of $\sigma\in S_{n}$. With this goal in mind, let $(a_{l})_{l=1}^{n}$ be the cycle type of $\sigma$. It is known that
$C_{S_{n}}(\sigma)\cong\prod_{l:a_{l}\neq 0}C_{l}^{a_{l}}\rtimes S_{a_{l}},$ (4)
where $C_{l}$ is the cyclic group of order $l$ and $S_{a_{l}}$ acts on $C_{l}^{a_{l}}$ by permuting the coordinates (see, for example [46, pg 295]). The bijection in equation (4) can be realized by noting that permutation in $C_{S_{n}}(\sigma)$ must cyclically shift and then permute the $l$-cycles of $\sigma$. This implies that the cycle type of $\tau\in C_{S_{n}}(\sigma)$ is the sum of the cycle-types of $\tau_{l}\in C_{l}^{a_{l}}\rtimes S_{a_{l}}$. Furthermore, the product in equation (4) implies that, under the uniform distribution on $C_{S_{n}}(\sigma)$, the permutations $(\tau_{l})_{l:a_{l}\neq 0}$ are independent. We have thus arrived at the following lemma.
Lemma 3.3.
Fix a permutation $\sigma$ with cycle type $(a_{l})_{l=1}^{n}$. Let $\tau$ be a uniform sample from $C_{S_{n}}(\sigma)$ and let $b=(b_{i})_{i=1}^{n}$ be the cycle type of $\tau$. Let $(\tau_{l})_{l:a_{l}\neq 0}$ be independent permutations with $\tau_{l}$ uniformly sampled from $C_{l}^{a_{l}}\rtimes S_{a_{l}}$. If $b_{l}=(b_{l,i})_{i=1}^{n}$ is the cycle type of $\tau_{l}$, then
$b\stackrel{{\scriptstyle d}}{{=}}\sum_{l:a_{l}\neq 0}b_{l}.$
Lemma 3.3 implies that to run the lumped Burnside process on $\mathscr{P}_{n}$ it is sufficient to know how to sample the cycle type of permutation $\tau$ uniformly sampled from $C_{l}^{a}\rtimes S_{a}\subseteq S_{la}$. The cycle type of uniformly sampled elements of $\Gamma^{a}\rtimes S_{a}$ for general $\Gamma\subseteq S_{l}$ was studied in [28]. The following proposition specializes some of their results to the case when $\Gamma=C_{l}$. For completeness, a proof is given in Appendix A.1.
Proposition 3.4.
Fix $l,a_{l}\geq 1$ and let $(\lambda_{j})_{j=1}^{m}$ be the cycle lengths of a uniformly distributed element of $S_{a_{l}}$. Let $(U_{j})_{j=1}^{m}$ be independent and uniformly distributed elements of $[l]$ and let $d_{j}=\gcd(U_{j},l)$. Let $(b_{i})_{i=1}^{la_{l}}$ be given by
$b_{i}=\sum_{j=1}^{m}d_{j}I[\lambda_{j}l/d_{j}=i],$
where $I[A]=1$ if $A$ is true and $I[A]=0$ otherwise. Then $(b_{i})_{i=1}^{la_{l}}$ has the same distribution as the cycle type of a uniformly sampled element of $C_{l}^{a_{l}}\rtimes S_{a_{l}}$.
The computational benefit of Proposition 3.4 is that the cycles lengths $(\lambda_{j})_{j=1}^{m}$ can be sampled directly without having to sample a permutation. In our implementation we sample $(\lambda_{j})_{j=1}^{m}$ via discrete stick breaking.
Definition 3.5.
The discrete stick breaking distribution on $n$ is a distribution over sequences $(\lambda_{j})_{j=1}^{m}$ with $\lambda_{j}\in\mathbb{Z}_{+}$ and $\sum_{j=1}^{m}\lambda_{j}=n$. To sample from the discrete breaking distribution first sample $\lambda_{1}$ uniformly from $[n]$. If $\lambda_{1}=0$, then stop. Otherwise, sample $\lambda_{2}$ uniformly from $[n-\lambda_{1}]$. If $\lambda_{1}+\lambda_{2}=n$, then stop and otherwise sample $\lambda_{3}$ uniformly from $[n-\lambda_{1}-\lambda_{2}]$. Continue in this way to produce $(\lambda_{j})_{j=1}^{m}$ with $\sum_{j=1}^{m}\lambda_{j}=n$.
It is well–known that the discrete stick breaking distribution is the same as the distribution of the cycle lengths of a randomly drawn partition $\sigma\in S_{n}$. The discrete stick breaking distribution is also an instance of Ewen’s distribution [18].
Together Lemma 3.3, Proposition 3.4 and Definition 3.5 give an efficient implementation of the lumped Burnside process. If we are currently at the partition $a=(a_{l})_{l\geq 1}$, then a single step is as follows
  • For each $l$ with $a_{l}\neq 0$ independently sample $b_{l}=(b_{l,i})_{i\geq 1}$ using Proposition 3.4 and return $b=\sum_{l:a_{l}\neq 0}b_{l}$.
A more detailed description is given in Algorithm 1. The following examples also give a feel for the procedure.
Algorithm 1 Lumped Burnside process for partitions
1:$(a_{l})_{l=1}^{n}$ (partition of $n$)
2:$(b_{l})_{l=1}^{n}$ (new partition sampled from the Burnside process)
3:$b_{l}\leftarrow 0$ for all $l\in[n]$
4:for $l\leftarrow 1$ to $n$ do
5:     if $a_{l}\neq 0$ then
6:         Sample $(\lambda_{j})_{j=1}^{m}$ from the discrete stick breaking distribution on $a_{l}$
7:         for $j\leftarrow 1$ to $m$ do
8:              Sample $U$ uniformly from $[l]$
9:              $d\leftarrow\gcd(U,l)$
10:              $i\leftarrow\lambda_{j}l/d$
11:              $b_{i}\leftarrow b_{i}+d$
12:         end for
13:     end if
14:end for
15:return $(b_{l})_{l=1}^{n}$
Example 3.6.
Suppose $a=(n,0,0,\ldots)$ so that $a_{l}\neq 0$ only for $l=1$. In this case, $(b_{i})_{i\geq 1}$ is simply the cycle type of a uniformly sample permutation in $S_{n}$. Equivalently, $b_{i}=\#\{j:\lambda_{j}=i\}$ where $(\lambda_{j})_{j=1}^{m}$ is drawn from the discrete stick breaking distribution on $n$.
Example 3.7.
Suppose $a=(0,0,\ldots,0,1)$ so that $a_{n}=1$ and $a_{l}=0$ otherwise. The new partition $(b_{i})_{i\geq 1}$ has exactly one non-zero entry $b_{n/d}=d$ where $d$ is a divisor of $n$. The probability of a particular value of $d$ is $\phi(n/d)/n$ where $\phi$ is Euler’s totient function. In other words, the partition $(b_{i})_{i\geq 1}$ is the cycle type of a uniformly chosen element of $C_{n}\subseteq S_{n}$.
The main benefit of Algorithm 1 is that we only have to store and use the non-zero values of $(a_{l})_{l=1}^{n}$. This can dramatically reduce the memory and time requirements of the algorithm. Indeed, the constraint $\sum_{1\geq 1}la_{l}=n$ implies that at most $2\sqrt{n}$ values of $a_{l}$ can be non-zero and that each $a_{l}$ is at most $n$. Thus, the memory required to store $(a_{l})_{l\geq 1}$ is $O(\sqrt{n}\log n)$.
The average time complexity of the lumped Burnside process is at most $O(\sqrt{n}(\log n)^{2})$. Sampling $(\lambda_{j})_{j=1}^{m}$ from the discrete stick breaking distribution on $a_{l}$ has average complexity $(\log a_{l})^{2}$. The complexity of sampling uniformly from $[l]$ and computing the greatest common divisor $d$ has complexity $O(\log l)$ and this has to be done $O(\log a_{l})$ times on average. Thus, the average case complexity of the algorithm is
$O\Bigg{(}\sum_{l:a_{l}\neq 0}(\log a_{l})^{2}+\log l\log a_{l}\Bigg{)}=O\Bigg{(}(\log n)^{2}|\{l:a_{l}\neq 0\}|\Bigg{)}=O\big{(}\sqrt{n}(\log n)^{2}\big{)}.$
In contrast, the original Burnside process on $S_{n}$ requires storing a permutation of size $n$ and so has linear memory requirements. Likewise, computing the cycle type of a permutation given as a list of numbers also has linear complexity. Thus, a single step of the original process has time and memory complexity of order $n\gg\sqrt{n}(\log n)^{2}$. These theoretical arguments are supported by our experiments which are reported in Figure 4.
However, while the per-step complexity of the lumped Burnside process is very low, we have found that it does not always mix quickly. Indeed, consider the process started a partition with a single part (i.e.$a_{n}=1$). Example 3.7 show that the chance that the chain remains at $a_{n}=1$ is $\phi(n)/n$. If $n$ is prime, this chance of holding is $1-1/n$ and the mixing time is bounded below by $n$. Figure 5(a) shows when the Burnside process is initialized at the all one’s partition (i.e. $a_{1}=n)$, the largest part is constant for long periods of time. The next section introduces a second Markov chain on $\mathscr{P}_{n}$ that addresses this concern.
Figure 4: The above is a log–log plot showing the benefits of the lumped process for partitions. Each point corresponds to running the corresponding Burnside process for 10,000 steps. The dashed lines have a fixed slopes (1 for the Burnside process, 1/2 for the lumped process and reflected process) and fitted intercepts. The good fit these lines supports our analysis that, up to log factors, the original process has complexity of order $n$ and the lumped process has complexity of order $\sqrt{n}$. Note also the comparable per-step complexity of the lumped and reflected processes.
3.3 Speeding up the Burnside process
The reflected Burnside process is a second Markov chain on $\mathscr{P}_{n}$ that empirically mixes much faster than the Burnside process. Importantly, the reflected Burnside process has the same per-step complexity as the lumped Burnside process from Section 3.2. Let $P$ be the transition kernel for the lumped Burnside process on $\mathscr{P}_{n}$ and let $\Pi:\mathscr{P}_{n}\times\mathscr{P}_{n}\to\{0,1\}$ be the permutation matrix with
$\Pi(a,b)=\begin{cases}1&\text{if }b\text{ is the transpose of }a,\\ 0&\text{otherwise}.\end{cases}$
The reflected Burnside process is the Markov chain with transition kernel $Q=\Pi P$. That is, to take one step of the reflected Burnside process, first compute the transpose of the current partition and then take a step of the lumped Burnside process via Algorithm 1. Since transposing is a bijection from $\mathscr{P}_{n}$ to $\mathscr{P}_{n}$, the transition kernel $Q$ is still ergodic with uniform stationary distribution. However, our simulations show that the transition matrix $Q$ mixes much more quickly than the original $P$. For example, Figure 5 shows that, under $Q$, the number of parts and the size of the largest part quickly approach the value predicted by Theorem 3.2. Under $P$, both features are far from their stationary distribution even after many steps.
(a) Largest part under the Burnside process $(n=10^{10})$.
(b) Number of parts under the Burnside process $(n=10^{10})$.
Figure 5: The above figures demonstrate the advantage of the reflected Burnside process over the lumped Burnside process on partitions of size $n=10^{10}$. . To generate these figures, both processes were initialized at the partition $1^{n}$ and run for 200 steps. Figure 5(a) shows the evolution of the size of the largest part and Figure 5(b) shows the evolution of the number of parts. Theorem 3.2 from [34] gives the asymptotic distribution of both these statistics under the uniform distribution and the asymptotic mean is represented as a dashed line. Under the reflected Burnside process, both statistics agree with the asymptotic distribution. Without transpositions, neither statistics is close to stationary. Figure 5(a) also shows that under the Burnside process there is a very strong autocorrelation in the size of the largest part. Indeed, if a partition contains a single part of size $l$, then, under the Burnside process, this part remains with probability $\phi(l)/l$. This autocorrelation is one barrier to mixing that the reflected Burnside process manages to avoid.
The findings in Figure 5 are consistent with a number of other simulations for different features and different values of $n$. For every feature that we have considered, the mixing of time of $Q$ seems to grow logarithmically in $n$ and remains less than 3030~{}3030 for $n$ as large as $10^{8}$.
An algorithm for computing the transpose of a partition in exponential notation is given in Algorithm 3 in the appendix. The complexity of Algorithm‘3 is on the order of $\sqrt{n}\log n$. This means that using $Q$ instead of $P$ increases the per-step complexity by at most a constant factor. The mixing time benefits of $Q$ far outweigh the moderate increase in per-step complexity.
The reflected Burnside process is an example of speeding up a Markov chain by adding a deterministic jump that preserves the stationary distribution. This technique is well–established in the literature [15, 14, 38, 12]. Proving a quantitative bound on the speed-up in this example is an open problem.
4 Double cosets
In this section we develop a general method for uniformly sampling double cosets via the Burnside process. This is specialized to contingency tables in Section 5. Through-out $H$ and $K$ will be subgroups of a finite group $G$. The group $H\times K$ acts on $G$ by
$s^{h,k}=h^{-1}sk.$ (5)
This gives an equivalence relation on $G$
$s\sim t\Longleftrightarrow t=h^{-1}sk\text{ for some }(h,k)\in H\times K.$
The equivalence classes are double cosets, $HsK$ denotes the double coset containing $s$ and $H\backslash G/K$ denotes the set of all double cosets. These definitions are standard in basic group theory [46, Page 23]. Classical facts are
$|HsK|$ $=\frac{|H||K|}{|H\cap sKs^{-1}|},$ (6)
$|H\backslash G/K|$ $=\frac{1}{|H||K|}\sum_{h,k}|G^{h,k}|,\quad G^{h,k}=\{s:s^{h,k}=s\}.$ (7)
Despite these nice formulas, enumerating or describing double cosets can be an intractable problem. [27] study and provide references for a host of examples where the enumeration problems are interesting. Many of the examples have connections to areas of mathematics beyond group theory. One example, contingency tables, is widely used in statistical applications and is developed in Section 5 below.
The Burnside process applies to the action of $H\times K$ on $G$ and generates uniformly chosen double cosets. In this case, the Burnside process is a Markov chain on $G$. To carry it out, two steps must be implemented:
  1. 1.
    From $s\in G$, uniformly choose $(h,k)$ in $(H\times K)_{s}=\{(h,k):s^{h,k}=s\}$.
  2. 2.
    From $(h,k)$, uniformly choose $t$ in $G^{h,k}=\{t:t^{h,k}=t\}$.
To implement these steps, here is a characterization of $(H\times K)_{s}$ and $G^{h,k}$. In the following, $C_{G}(h)$ denotes the centralizer of $h$ in $G$. That is $C_{G}(h)=\{s\in G:sh=hs\}$.
Lemma 4.1.
Let $H\times K$ act on $G$ as in (5). Then
  1. 1.
    For all $s\in G,h\in H$ and $k\in K$, $s^{h,k}=s$ if and only if $k=s^{-1}hs$.
  2. 2.
    For all $s\in G$, $(H\times K)_{s}=\{(h,k):h\in H\cap sKs^{-1}\text{ and }k=s^{-1}hs\}$.
  3. 3.
    For all $(h,k)\in H\times K$, if $G^{h,k}$ is non-empty, then $G^{h,k}=C_{G}(h)s$ where $s$ is any fixed element of $G^{h,k}$.
Proof.
The first claim is simply stating that $h^{-1}sk=s$ if and only if $k=s^{-1}hs$. The first claim implies that if $(h,k)\in(H\times K)_{s}$, then $k$ is determined by the relation $k=s^{-1}hs$. The set of possible $h$ is
$H\cap\{h:s^{-1}hs\in K\}=H\cap sKs^{-1}.$
This implies that $h\mapsto(h,s^{-1}hs)$ is a bijection between $H\cap sKs^{-1}$ and $(H\times K)_{s}$ which is the second claim. For the final claim, suppose that there exists some $s\in G^{h,k}$. Then $hs=sk$ and furthermore $t\in G^{h,k}$ if and only if $ht=tk$. Thus, $t\in G^{h,k}$ is equivalent to $h(ts^{-1})=(ts^{-1})h$. This implies that $G^{h,k}=C_{G}(h)s$. ∎
Lemma 4.1 gives the following description of the Burnside process for transitioning from $s$ to $t$ in this example.
  1. 1.
    From $s\in G$, choose $h$ uniformly from $H\cap sKs^{-1}$.
  2. 2.
    From $h$, choose $g$ uniformly in $C_{G}(h)$ and set $t=gs$.
Thus, to run the Burnside process of double cosets it suffices to be able to sample from the subgroups $H\cap sKs^{-1}$ and $C_{G}(h)$ for all $s\in G$ and $h\in H$.
4.1 Graph coloring
In [11], double cosets are used to enumerate graph colorings. Let $\mathscr{G}$ be a graph with vertex set $[n]$. We will let $K\subseteq S_{n}$ be the automorphism group of $\mathscr{G}$. That is $K$ consists of the permutations $\sigma$ that preserve the edges in $\mathscr{G}$. Also let $\lambda=(\lambda_{1},\ldots,\lambda_{I})$ be a partition of $\lambda$ so that $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{I}=n$. We wish to color the vertices of $\mathscr{G}$ with $I$ colors so that $\lambda_{1}$ vertices get the first color, $\lambda_{2}$ vertices get the second and so on. The goal of [11] is to count the number of such colorings modulo the symmetries in the automorphism group $K$.
To see the connection with double cosets, let $S_{\lambda}\subseteq S_{n}$ be the subgroup of permutations that permute $1,\ldots,\lambda_{1}$ amongst themselves, $\lambda_{1}+1,\ldots,\lambda_{1}+\lambda_{2}$ amongst themselves and so on. The subgroup $S_{\lambda}$ is a parabolic subgroup, see Section 5.1 for more details. Given $\sigma\in S_{n}$, we can construct a coloring of $\mathscr{G}$ be assigning a vertex $v$ the $i$th color if $\sigma(v)>\lambda_{1}+\cdots+\lambda_{i-1}$ and $\sigma(v)\leq\lambda_{1}+\cdots+\lambda_{i}$. Two permutations $\sigma$ and $\tau$ give isomorphic colorings if and only if $S_{\lambda}\sigma K=S_{\lambda}\tau K$. Thus, distinct graph colorings are in bijection with the $S_{\lambda}\backslash S_{n}/K$ double cosets. [11] use the lexicographical ordering on $S_{n}$ to recursively enumerate these double cosets and hence graph colorings. For larger problems, the Burnside process can be used to sample graph colorings.
5 Contingency tables
A contingency table is an $I\times J$ array of non-negative integers $(T_{i,j})$ frequently used to record cross classified data. For example, Tables 1 and 2 show two classical data sets; the first is a classification by hair color and eye color. The second is a classification by number of children and yearly income. A standard task of statistical analysis of such cross classified data is to test for independence of the two variables. This test is most often done by computing the ‘chi-square statistic’
$\chi^{2}=\chi^{2}(T)=\sum_{i,j}\frac{(T_{i,j}-r_{i,\cdot}c_{\cdot,j}/n)^{2}}{r_{i,\cdot}c_{\cdot,j}/n},$ (8)
with $r_{i,\cdot}=\sum_{j}T_{i,j}$, $c_{\cdot,j}=\sum_{i}T_{i,j}$ and $n=\sum_{i,j}T_{i,j}$. Standard practice compares the chi-square statistic to its limit distribution–the ‘chi-squared distribution on $(I-1)(J-1)$ degrees of freedom’. To illustrate, in Table 1, $I=J=4$, $\chi^{2}=138.29$ on $3\times 3=9$ degrees of freedom. The value $138.29$ is way out in the tail of the limiting distribution and the standard test rejects the null hypothesis of independence.
Table 1: Eye Color vs hair color for $n=592$ subjects. Reproduced from Table 1 in [19]
Hair Color
Eye Color Black Brunette Red Blond Total
Brown 68 119 26 7 220
Blue 20 84 17 94 215
Hazel 15 54 14 10 93
Green 5 29 14 16 64
Total 108 286 71 127 592
Table 2: Number of children vs yearly income for $n=25,263$ Swedish families. Reproduced from Table 2 in [19]
Yearly Income
Units of 1,000 Kroner
Number of Children 0–1 1–2 2–3 3+ Total
0 2161 3577 2184 1636 9558
1 2755 5081 $2$ 1052 11110
2 936 1753 640 306 3635
3 225 419 96 38 778
$\geq 4$ 39 98 31 14 182
Total 6116 10928 5173 3046 25263
The statistical background is covered in any introductory statistical text and [4] is an extensive survey. Early on, R.A. Fisher [33] suggested looking at the conditional distribution of the chi-square statistic when the row and column sums are fixed and the two variables are independent. In a later development, Diaconis and Efron [19] suggested looking at the uniform distribution on tables with fixed row and columns sums as a way of interpreting ‘large’ values of $\chi^{2}$.
[19] led to a healthy development, now subsumed under the label ‘algebraic statistics’. The main questions are can one enumerate all such tables or pick one at random. These problems turn out to be #$\#P$-complete, meaning exact enumeration is impossible and a host of approximations have been developed. An influential paper of Diaconis and Sturmfels [26] develops Markov chain Monte Carlo approximations. It also contains a comprehensive review of conditional testing. See [5] for recent references.
This section develops an algorithm for uniform generation of contingency tables with given row and column sums. The algorithm uses the Burnside process for the $S_{\lambda},S_{\mu}$ double cosets in the symmetric group $S_{n}$ and previously appeared in [23, Chapter 5]. Here $S_{\lambda}$ and $S_{\mu}$ are ‘Young subgroups’ corresponding to the row sums $\lambda$ and column sums $\mu$. Section 5.1 explains the connection between contingency tables and double cosets and sets up notation. Section 5.2 describes the lumped Burnside process for contingency tables. This lumped process is applied to Tables 1 and 2 in Section 5.3. We compare the results with previous efforts [19, 26] and summarize our experiences with our examples.
5.1 Contingency tables as double cosets
This section explains the relationship between contingency tables with given row and column sums and double cosets of $S_{n}$. We roughly follow Section 1.3 of [40]. Through-out $\lambda=(\lambda_{i})_{i=1}^{I}$ and $\mu=(\mu_{j})_{j=1}^{J}$ are two compositions of $n$. That is, $\lambda_{i}$ and $\mu_{j}$ are positive integers with $\lambda_{1}+\cdots+\lambda_{I}=\mu_{1}+\cdots+\mu_{J}=n$. Since $\lambda$ and $\mu$ are compositions of $n$ rather than partitions, we do not assume that they are in non-decreasing order.
The space of contingency tables with rows sums $\lambda$ and column sums $\mu$ will be denoted by $\mathscr{T}_{\lambda,\mu}$. That is
$\mathscr{T}_{\lambda,\mu}=\left\{T\in\mathbb{Z}^{I\times J}_{\geq 0}:\sum_{j=1}^{J}T_{i,j}=\lambda_{i}\text{ for all $i$ and }\sum_{i=1}^{I}T_{i,j}=\mu_{j}\text{ for all }j\right\}.$ (9)
The composition $\lambda$ defines a set partition $L=(L_{i})_{i=1}^{I}$ with
$L_{1}=\{1,\ldots,\lambda_{1}\},L_{2}=\{\lambda_{1}+1,\ldots,\lambda_{1}+\lambda_{2}\},\ldots,L_{I}=\{n-\lambda_{I}+1,\ldots,n\}.$
Likewise, $\mu$ defines the set partition $(M_{j})_{j=1}^{J}$ with $M_{1}=\{1,\ldots,\mu_{1}\}$,$M_{2}=\{\mu_{1}+1,\ldots,\mu_{1}+\mu_{2}\}$,…,$M_{J}=\{n-\mu_{J}+1,\ldots,n\}$.
For any set partition $(A_{i})_{i=1}^{I}$, define a subgroup $S_{A}\subseteq S_{n}$ by
$S_{A}=\{\sigma\in S_{n}:\sigma(A_{i})=A_{i}\text{ for all }i\}.$ (10)
Permutations in $S_{A}$ permute the elements of $A_{1}$ amongst themselves, the elements of $A_{2}$ amongst themselves and so on. Thus, $S_{A}\cong\prod_{i=1}^{I}S_{|A_{i}|}$. When the blocks of the set partition are contiguous, as is the case with $(L_{i})_{i=1}^{I}$ and $(M_{j})_{j=1}^{J}$, the subgroup $S_{A}$ is called a Young subgroup or a parabolic subgroup. For the compositions $\lambda$ and $\mu$, we will use $S_{\lambda}$ and $S_{\mu}$ to denote $S_{L}$ and $S_{M}$. The following map will be used to define a bijection between $S_{\lambda}\backslash S_{n}/S_{\mu}$ and $\mathscr{T}_{\lambda,\mu}$.
Definition 5.1.
Let $(L_{i})_{i=1}^{I}$ and $(M_{j})_{j=1}^{J}$ be the set partitions above. Define a function $f:S_{n}\to\mathbb{Z}_{\geq 0}^{I\times J}$ by
$f(\sigma)_{i,j}=|L_{i}\cap\sigma(M_{j})|.$
The image of $f$ is the space of double cosets $\mathscr{T}_{\lambda,\mu}$ and it can be shown that $f(\sigma)=f(\tau)$ if and only if $S_{\lambda}\sigma S_{\mu}=S_{\lambda}\tau S_{\mu}$. In particular, we have the following result.
Proposition 5.2 (Theorem 1.3.10 in [40]).
The function $f$ in Definition 5.1 is constant on the double costs in $S_{\lambda}\backslash S_{n}/S_{\mu}$ and induces a bijection between $S_{\lambda}\backslash S_{n}/S_{\mu}$ and $\mathscr{T}_{\lambda,\mu}$.
Sampling $\sigma\in S_{n}$ uniformly and then computing $f(\sigma)$ induces a probability measure on $\mathscr{T}_{\lambda,\mu}$ called the Fisher—Yates distribution. Under the Fisher–Yates distribution, the probability of a contingency table $T=f(\sigma)$ is proportional to $|S_{\lambda}\sigma S_{\mu}|$ which can be computed using (6). Note that
$S_{\lambda}\cap\sigma S_{\mu}\sigma^{-1}=S_{L}\cap S_{\sigma M}=S_{L\land\sigma M},$
where $\sigma M$ is the set partition $(\sigma(M_{j}))_{j=1}^{J}$ and $L\land\sigma M$ is the meet of $L$ and $\sigma M$ meaning
$L\land\sigma M=(L_{i}\cap\sigma(M_{j}))_{i=1,j=1}^{I,J}.$
The subgroup $S_{\lambda}\cap\sigma S_{\mu}\sigma^{-1}$ therefore has size $\prod_{i,j}f(\sigma)_{i,j}!$. These observations imply the following.
Proposition 5.3.
Let $p$ be the probability mass function for the Fisher–Yates distribution on $\mathscr{T}_{\lambda,\mu}$, then for all $T\in\mathscr{T}_{\lambda,\mu}$,
$p(T)=\frac{1}{n!}\frac{\left(\prod_{i=1}^{I}\lambda_{i}!\right)\left(\prod_{j=1}^{J}\mu_{j}!\right)}{\prod_{i,j}T_{i,j}!}.$ (11)
5.2 The Burnside process for contingency tables
In this section we explain the lumped Burnside process on contingency tables. That is, we describe the transitions from $f(\sigma)\in\mathscr{T}_{\lambda,\mu}$ to $f(\tau)$ when $\sigma\in S_{n}$ transitions to $\tau$ according to the Burnside process for $S_{\lambda},S_{\mu}$ double cosets. Later, we will quantify the computational benefits of the lumped process over the original process. These benefits are also demonstrated empirically in the next section using the two examples in Tables 1 and 2. A similar description of the lumped Burnside process has previously appeared in [23]. However, their implementation has complexity that scales linearly in $n$, the sum of the table entries. As shown later, the average case complexity of our algorithm is on the order of $(\log n)^{2}$.
The following lemma contains the key insights behind our implementation of the lumped Burnside process. Specifically, it describes the distribution of $f(\tau)\in\mathscr{T}_{\lambda,\mu}$ when $\tau$ is drawn uniformly from the set $S_{n}^{h,k}$ as in the second step of the Burnside process. The definitions of $\lambda,\mu,(L_{i})_{i=1}^{I},(M_{j})_{j=1}^{J}$ and $f$ are the same as in the previous section. As before $S_{\lambda}$ and $S_{\mu}$ are Young subgroups. Elements $h\in S_{\lambda}$ and $k\in S_{\mu}$ can be represented as $h=(h_{i})_{i=1}^{I}$ and $k=(k_{j})_{j=1}^{J}$ where $h_{i}$ and $k_{j}$ are permutations of $L_{i}$ and $M_{j}$ respectively. For any permutation, $h$, we will let $\mathscr{C}(h)$ denote the set of cycles of $h$.
Lemma 5.4.
Let $h=(h_{i})_{i=1}^{I}\in S_{\lambda}$ and $k=(k_{j})_{j=1}^{J}\in S_{\mu}$ be such that $S_{n}^{h,k}=\{\sigma\in S_{n}:h^{-1}\sigma k=\sigma\}\neq\emptyset$. Let $\tau$ be drawn uniformly from $S_{n}^{h,k}$, and, for $l\in[n]$, define $X^{(l)}\in\mathbb{Z}_{\geq 0}^{I\times J}$ by
$X^{(l)}_{i,j}=|\{C:|C|=l,C\in\mathscr{C}(h_{i})\text{ and }\tau^{-1}(C)\in\mathscr{C}(k_{j})\}|.$
Then, $X^{(l)}$ is distributed according to the Fisher–Yates distribution with row sums $r^{(l)}_{i}=|\{C\in\mathscr{C}(h_{i}):|C|=l\}|$ and column sums $c^{(l)}_{j}=|\{C\in\mathscr{C}(k_{j}):|C|=l\}|$. Furthermore,
$f(\tau)=\sum_{l=1}^{n}lX^{(l)}.$
Lemma 5.4 is proved in Appendix A.2. To get a feel for the lemma, the following examples are helpful.
Example 5.5.
Suppose that $h$ and $k$ are both the identity permutation. In this case $r^{(1)}_{i}=\lambda_{i}$, $c^{(1)}_{j}=\mu_{j}$ and $r^{(l)}_{i}=c^{(l)}_{j}=0$ otherwise. We also have that $S_{n}^{h,k}=S_{n}$. Thus, Lemma 5.4 is simply stating that if $\tau$ is uniformly distributed in $S_{n}$, then $f(\tau)$ is distributed according to the Fisher–Yates distribution.
Example 5.6.
Suppose that $\lambda=\mu=(m,\ldots,m)$ ($I$ times) and that for each $i$, $h_{i}=k_{i}$ is an $m$-cycle. Then $r^{(m)}_{i}=c^{(m)}_{i}=1$ and $r^{(l)}_{i}=r^{(l)}_{j}=0$ for $l\neq m$. In this case Lemma 5.4 is stating that $f(\tau)=mX^{(m)}$ and that $X^{(m)}\in\mathbb{Z}_{\geq 0}^{I\times I}$ is a uniformly drawn permutation matrix.
Lemma 5.4 has the following corollary which is useful for the lumped Burnside process.
Corollary 5.7.
Let $h$, $k$ and $\tau$ be as in Lemma 5.4. The distribution of $f(\tau)$ only depends on the vectors $(r^{(l)})_{l\geq 1}$ and $(c^{(l)})_{l\geq 1}$ which are the cycle types of $(h_{i})_{i=1}^{I}$ and $(k_{j})_{j=1}^{J}$ respectively.
Corollary 5.7 states that, when running the lumped Burnside process, we do not need to sample the full permutations $h$ and $k$. It is sufficient to just sample $r_{i}^{(l)}$ and $c_{j}^{(l)}$ from the correct distribution. The following lemma gives describes the distribution of $r_{i}^{(l)}$ and $c_{j}^{(l)}$.
Lemma 5.8.
Fix $T\in\mathscr{T}_{\lambda,\mu}$ and let $\sigma$ be any permutation with $f(\sigma)=T$. Let $(h,k)$ be uniformly sampled from $(S_{\lambda}\times S_{\mu})_{\sigma}$. Define $a_{i,j}^{(l)}$ to be the number of $l$-cycles of $h$ contained in $L_{i}\cap\sigma(M_{j})$ and let $a_{i,j}=(a_{i,j}^{(l)})_{l\geq 1}$. Then, $(a_{i,j})_{i,j}$ are independent and $a_{i,j}$ is equal in distribution to the cycle type of a permutation of length $T_{i,j}$. Furthermore, for every $l$
$r^{(l)}_{i}=\sum_{j=1}^{J}a_{i,j}^{(l)}\quad\text{and}\quad c^{(l)}_{j}=\sum_{i=1}^{I}a_{i,j}^{(l)},$
where $r^{(l)}_{i}=|\{C\in\mathscr{C}(h):C\subseteq L_{i},|C|=l\}|$ and $c^{(l)}_{j}=|\{C\in\mathscr{C}(k):C\subseteq M_{j},|C|=l\}|$.
The proof of Lemma 5.8 is given in Appendix A.3. In our implementation, $(a_{i,j}^{(l)})$ are sampled using discrete stick breaking as described in Definition 3.5. The two steps of the Burnside process for contingency tables are
  1. 1.
    From a contingency table $T$, sample $(a_{i,j}^{(l)})$ using discrete stick breaking and compute $(r^{(l)})_{l\geq 1}$ and $(c^{(l)})_{l\geq 1}$ as in Lemma 5.8.
  2. 2.
    Given $(r^{(l)})_{l\geq 1}$ and $(c^{(l)})_{l\geq 1}$ sample $X^{(l)}$ and return $T^{\prime}=\sum_{l\geq 1}lX^{(l)}$ as in Lemma 5.4.
A full description of the lumped chain is given at the end of this section in Algorithm 2. The lumped chain has both memory and speed advantages over the original process. The memory advantages are easier to see. To run the original process, we need to store the permutation $n$ with has memory requirement of the order $n\log n$. For the lumped process, we need to store the table $T$ which requires on the order of $IJ\log n$ memory. The lumped chain also requires storing the non-zero values of $r_{i}^{(l)}$ and $c_{j}^{(l)}$. The number of non-zero values of $r_{i}^{(l)}$ and $c_{j}^{(l)}$ is upper bounded by the number of non-zero values of $a_{i,j}^{(l)}$. Since $a_{i,j}^{(l)}$ is drawn from a discrete stick breaking distribution, there are on average $O(\log T_{i,j})$ values of $l$ with $a_{i,j}^{(l)}\neq 0$. Thus, the total memory requirement of the lumped chain is $O(IJ\log n)$ on average.
The speed advantages of the lumped chain are also substantial. The values $(a_{i,j}^{(l)})_{l\geq 0}$ can be sampled using stick breaking with an average case complexity on the order of $(\log T_{i,j})^{2}$. Therefore, $(r^{(l)})_{l\geq 1}$ and $(c^{(l)})_{l\geq 1}$ can all be computed with complexity $O(IJ(\log n)^{2})$. Given $(r^{(l)})_{l\geq 1}$ and $(c^{(l)})_{l\geq 1}$, we only need to sample $X^{(l)}$ for the non-zero values of $r^{(l)}$. Thus, for all but $IJ\log n$ values of $l$, $X^{(l)}$ is zero and does not need to be computed. For the non-zero values of $r^{(l)}$, $X^{(l)}$ can be sampled with complexity $O(IJ\log n)$. Thus, the time complexity of the lumped chain is $O((IJ\log n)^{2})$. This is a substantial speed up over the original chain which has linear complexity. Indeed, simply computing $f(\sigma)$ has complexity of order $n$.
Algorithm 2 Lumped Burnside process for contingency tables
1:$T$ (contingency table of size $I\times J$)
2:$T^{\prime}$ (new contingency table sampled from the Burnside process) $\triangleright$ Step 1
3:$r_{i}^{(l)}\leftarrow 0$ for $i\in[I]$ and $l\in[n]$
4:$c_{j}^{(l)}\leftarrow 0$ for $j\in[J]$ and $l\in[n]$
5:for $i\leftarrow 1$ to $I$ do
6:     for $j\leftarrow 1$ to $J$ do
7:         Sample $(\lambda_{m})_{m=1}^{M}$ from the discrete stick breaking distribution on $T_{i,j}$
8:         for $m\leftarrow 1$ to $M$ do
9:              $r_{i}^{(\lambda_{m})}\leftarrow r_{i}^{(\lambda_{m})}+1$
10:              $c_{j}^{(\lambda_{m})}\leftarrow c_{j}^{(\lambda_{m})}+1$
11:         end for
12:     end for
13:end $\triangleright$ Step 2
14:$T^{\prime}_{i,j}\leftarrow 0$ for $i\in[I]$ and $j\in[J]$
15:for $l$ such that $r^{(l)}\neq 0$ do
16:     Sample $X^{(l)}$ from the Fisher–Yates distribution with row sums $r^{(l)}$ and column sums $c^{(l)}$
17:     $T^{\prime}\leftarrow T^{\prime}+lX^{(l)}$
18:end for
19:return $T^{\prime}$
5.3 Applications of the Burnside process for contingency tables
In this section, we run the Burnside process on the spaces of contingency tables corresponding to Tables 1 and 2. Our simulations agree with the results reported in [19] and highlight the benefit of the lumped process.
Both Tables 1 and 2 have values of the chi-square statistic (8) that are very large compared to the chi-square statistics of tables drawn from the Fisher–Yates distribution (11). In such settings, [19] suggest comparing the chi-square statistic to the statistics of tables drawn from the uniform distribution on $\mathscr{T}(\lambda,\mu)$. Specifically, they suggest computing the ‘volume test statistic’ given by
$V(T)=\frac{\left|\{T^{\prime}\in\mathscr{T}(\lambda,\mu):\chi^{2}(T^{\prime})\geq\chi^{2}(T)\}\right|}{\left|\mathscr{T}(\lambda,\mu)\right|}.$
That is, $V(T)$ is the proportion of tables $T^{\prime}\in\mathscr{T}(\lambda,\mu)$ that have a larger chi-square statistic than $T$. Here $\lambda$ and $\mu$ are the row and column sums of $T$. Large values of $V(T)$ are evidence of stronger dependence between the row and column variables in $T$. Note that $V(T)$ can also be written as
$V(T)=\mathbb{P}(\chi^{2}(T^{\prime})\geq\chi^{2}(T)),$ (12)
where $T^{\prime}$ is drawn from the uniform distribution on $\mathscr{T}(\lambda,\mu)$. Thus, $V(T)$ can be approximated by running the Burnside process on $\mathscr{T}(\lambda,\mu)$ and reporting the proportion of sampled tables with $\chi^{2}(T^{\prime})\geq\chi^{2}(T)$. The results of such a simulation are reported in Tables 3. The simulation agrees with the findings in [19]. They find that Table 2 has a value of $\chi^{2}(T)$ that is much smaller than expected under the uniform distribution. On the other hand, Table 1 has a value of $\chi^{2}(T)$ that is fairly typical under the uniform distribution. The numerical value of $V(T)$ for Table 1 agrees with the value reported in Section 6 of [21]. They use a different Markov chain to estimate $V(T)$ and report that $V(T)\approx 0.154$.
Table 3: Estimates of the volume test statistic $V(T)$ for Tables 1 and 2. Each estimate was produced by running the lumped Burnside process for $2\times 10^{6}$ steps. The Burnside process was initialized at either Table 1 or Table 2 and run for $10^{4}$ steps as a “burn-in”. For each table, this procedure is repeated five times to get five estimates of $V(T)$. These five estimates and their median are reported for each table. On a personal computer, 1 hour was needed to perform all $10^{7}+5\times 10^{4}$ steps for Table 1. For Table 2, 4 hours and 20 minutes were needed to perform all $10^{7}+5\times 10^{4}$ steps.
Run 1 Run 2 Run 3 Run 4 Run 5 Median
Table 1 0.1545 0.1534 0.1532 0.1535 0.1533 0.1534
Table 2 $1.6\times 10^{-5}$ $1.35\times 10^{-5}$ $6.0\times 10^{-6}$ $8.0\times 10^{-6}$ $1.35\times 10^{-5}$ $1.35\times 10^{-5}$
There are many methods that use Markov chain Monte Carlo in the context of hypothesis testing to estimate a tail probability such as (12). The method we use to estimate $V(T)$ in Table 3 is consistent meaning that the estimates converge to $V(T)$ as the number of samples goes to infinity. However, the method does not have the error control guarantee that is sometimes desired for hypothesis testing. The methods in [8, 17, 16, 37] all provide such a guarantee. The remarkable thing about these methods is that they assume nothing about the mixing time of the Markov chain. Any of these methods can be used in conjunction with the lumped Burnside process.
Finally, Figure 6 shows the computational benefits of the lumped Burnside process. This figure, and the analysis in Section 5.2, show that the lumped process gives an exponential speed up over the unlumped process. The unlumped process has complexity that is linear in $n$ (the sum of the table entries). However, the complexity of the lumped process grows like a power of $\log n$.
Figure 6: The above figure is a linear–log plot showing the per-step complexity of the Burnside process for contingency tables. The plot was generated by running the Burnside process for 10,000 steps on scaled copies of Tables 1 and 2. Straight line approximations are shown as dashed lines. We see that if the entries of the table are increased by a multiplicative factor, the run time increases by additive factor. This supports our analysis which shows that the lumped process has a run time that is at most $O((\log n)^{2})$ for tables with fixed dimensions.
6 Conclusion
Classifying the orbits of a finite group acting on a finite set is a huge, unmanageable topic (!). The survey [42] points to the richness and depth within group theory. Our examples show that there are worthwhile applications further afield in statistics and probability.
The analysis and simulations in Sections 5 and 3 demonstrate the computational benefits of the lumped Burnside process. Even for partitions or contingency tables of moderate size, the lumped chain is several orders of magnitude faster than original process. We have found two common ideas that help achieve the full benefits of lumping. These are using simple data structures for the orbits, and using discrete stick breaking to generate the cycle type of a random permutation. We believe that both of these will help in other examples, and we give some more details below.
In both Sections 3 and 5 the orbits of the group action could be easily identified. In Section 3 the orbits were in bijection with partitions and in Section 5 the orbits were in bijection with contingency tables. These bijections meant we could run the Burnside process directly on the orbits. This led to an implementation that required far less memory than the original Burnside process. We also found that the choice of how to represent the orbits matters. In Section 3 we represented the partition in exponential notation and found that this worked better than the representing the partition as a list of parts. Unfortunately, in many other examples there is not a known identification of the orbits. Thus, this aspect of lumping may not be implementable in other examples.
The second aspect of lumping that we found helpful is discrete stick breaking. Lemmas 3.3 and 5.8 describe the distribution of the cycle type of a uniformly distributed element of a stabilizer of the group action. These results are useful because in both cases the second step of the Burnside process only depends on these cycles and not on the full permutation. We believe that this should hold in some generality. Specifically, when the group $G$ is a subgroup of a permutation group, then some form of stick breaking could be used to generate the cycle type of a uniformly chosen element of a stabilizer. This technique could be applied in cases when the orbits do not have a known representation.
For partitions, we also found that reflected Burnside process in Section 3.3 mixes much more quickly than the original Burnside process. The transpose map used by the reflected Burnside process is a bijection on partitions. However, the transpose map does not lift to a bijection on permutations. Thus, another benefit of lumping is that it can lead to new Markov chains that combine the Burnside process with deterministic bijections.
Finally, we end with two comments about implementation. First, the choice of programming language matters. We originally implemented our algorithms in Python but later moved to Julia. In Julia, the programs ran 100 times faster and could scale to larger problems. Second, both of our algorithms naturally lend themselves to parallelization. We did not explore this in our implementation, but it could lead to further speed–ups.
Acknowledgements
We thank Timothy Sudijono, Nathan Tung, Chenyang Zhong, Andrew Lin, Balazs Gerncser and Laurent Bartholdi for helpful comments and suggestions.
Appendix A Proofs
A.1 Proof of Proposition 3.4
Proof.
Let $\tau=(V_{1},V_{2},\ldots,V_{a_{l}};\pi)$ be a uniformly sampled element of $C_{l}^{a_{l}}\rtimes S_{a_{l}}$. Thus, $\pi$ is uniformly distributed in $S_{a_{l}}$, $V_{1},\ldots,V_{a_{l}}$ are uniformly distributed on $C_{l}$ and $(V_{1},\ldots,V_{a_{l}};\pi)$ are mutually independent. As an element of $S_{la_{l}}$, the permutation $\tau$ cyclically shifts the elements of $L_{1}=\{1,\ldots,l\}$ by $V_{1}$, the elements of $L_{2}=\{l+1,\ldots,2l\}$ by $V_{2}$, …, the elements of $L_{a_{l}}=\{(a_{l}-1)l+1,\ldots,a_{l}l\}$ by $V_{a_{l}}$ and then permutes the blocks $\{L_{1},\ldots,L_{a_{l}}\}$ according to $\pi$. Thus, if $i=kl+r$ with $r\in\{1,\ldots,l\}$, then $\tau(i)=\pi(k)l+(r+V_{k}\bmod l)$. Our goal is to understand the cycles of $\tau$.
Let $B_{1},\ldots,B_{m}$ be the cycles of $\pi$ and let $\lambda_{j}=|B_{j}|$ be the length of the $j$th cycle and write $B_{j}=(k_{1},k_{2},\ldots,k_{\lambda_{j}})$. The permutation $\tau$ thus cycles the blocks $L_{k_{1}},L_{k_{2}},\ldots,L_{k_{\lambda_{j}}}$. When applying $\tau$ $\lambda_{j}$ times, an element $i\in L_{k_{1}}$ returns to the block $L_{k_{1}}$ but will be shifted by $U_{j}=V_{k_{1}}+V_{k_{2}}+\cdots+V_{k_{\lambda_{j}}}$. Thus, the size of the cycle of $\tau$ that contains $i$ is $\lambda_{j}$ times the order of $U_{j}$ in $C_{l}$. The order of $U_{j}$ is $l/\gcd(U_{j},l)$ and thus each $i\in L_{k_{1}}$ is in a cycle of size $\lambda_{j}l/d_{j}$ where $d_{j}=\gcd(U_{j},l)$. Furthermore, the elements of $L_{k_{1}}\cup L_{k_{2}}\cup\cdots\cup L_{k_{\lambda_{j}}}$ are partitioned into $d_{j}$ cycles each of size $\lambda_{j}l/d_{j}$. Finally, since each $V_{k}$ is uniformly distributed in $\mathbb{Z}_{l}$, $U_{j}$ is also uniformly distributed. Thus, as stated in Proposition 3.4, every cycle length $\lambda_{j}$ contributes $d_{j}$ cycles of size $\lambda_{j}l/d_{j}$ to the cycle type of $\tau$. ∎
A.2 Proof of Lemma 5.4
Proof.
Let $\tau$ be drawn uniformly from $S_{n}^{h,k}\neq\emptyset$. Then, by Lemma 4.1, $\tau$ is drawn uniformly from the set $\{\tau\in S_{n}:k=\tau^{-1}h\tau\}$. In particular, the condition $S_{n}^{h,k}\neq\emptyset$ implies that $h=(h_{i})_{i=1}^{I}\in S_{\lambda}$ and $k=(k_{j})_{j=1}^{I}\in S_{\mu}$ are conjugate in $S_{n}$. Furthermore, for every $l$, the permutation $\tau^{-1}$ induces a bijection from the $l$-cycles of $k$ to the $l$-cycles of $h$. Since $\tau$ is uniformly distributed the bijection on the $l$-cycles is also uniformly distributed. The value $X_{i,j}^{(l)}$ is the number of $l$ cycles of $h$ in $L_{i}$ that $\tau^{-1}$ maps to an $l$ cycle of $k$ in $M_{j}$. Since the bijection on the $l$ cycles is uniformly distributed, $X^{(l)}$ has the Fisher–Yates distribution with row sums $r^{(l)}$ and column sums $c^{(l)}$.
For the last claim, we have
$f(\tau)_{i,j}$ $=|L_{i}\cap\tau(M_{j})|$
$=\sum_{l\geq 1}|\{x:x\in C,|C|=l,C\in\mathscr{C}(h_{i}),\tau^{-1}(C)\in\mathscr{C}(k_{j})\}|$
$=\sum_{l\geq 1}l\{C:|C|=l,C\in\mathscr{C}(h_{i}),\tau^{-1}(C)\in\mathscr{C}(k_{j})\}|$
$=\sum_{l\geq 1}lX_{i,j}^{(l)},$
as required. ∎
A.3 Proof of Lemma 5.8
Proof.
Let $(h,k)$ be drawn uniformly from $(S_{\lambda}\times S_{\mu})_{\sigma}$ where $\sigma$ is such that $f(\sigma)=T$. By Lemma 4.1, we know that $h$ is uniformly drawn from $S_{\lambda}\cap\sigma S_{\mu}\sigma^{-1}=S_{L\land\sigma(M)}$. It follows that $h=(h_{i,j})_{i,j}$ where $h_{i,j}$ is a uniformly drawn permutation of $L_{i}\cap\sigma(M_{j})$ and $(h_{i,j})_{i,j}$ are independent. The vector $(a_{i,j}^{(l)})_{l\geq 0}$ is simply the cycle type of $h_{i,j}$. Since $|L_{i}\cap\sigma(M_{j})|=T_{i,j}$, the vector $(a_{i,j}^{(l)})_{l\geq 0}$ has the discrete stick breaking distribution on $T_{i,j}$. Independence of $(a_{i,j})_{i,j}$ follows from the independence of $(h_{i,j})_{i,j}$.
Furthermore, since $h\in S_{\lambda}\cap\sigma S_{\mu}\sigma^{-1}$, every cycle of $h$ must be contained in $L_{i}\cap\sigma(M_{j})$ for some $i$ and $j$. Thus,
$r_{i}^{(l)}=|\{C\in\mathscr{C}(h_{i}):|C|=l,C\subseteq L_{i}\}|=\sum_{j=1}^{J}|\{C\in\mathscr{C}(h_{i}):|C|=l,C\subseteq L_{i}\cap\sigma(M_{j})\}|=\sum_{j=1}^{J}a_{i,j}^{(l)}.$
Finally, by Lemma 4.1 $k=\sigma^{-1}h\sigma$. This implies that if $C$ is an $l$-cycle of $h$ in $L_{i}\cap\sigma(M_{j})$, then $\sigma^{-1}(C)$ is an $l$-cycle of $k$ in $\sigma(L_{i})\cap M_{j}$. And so, by similar reasoning, $c^{(l)}_{j}=\sum_{i=1}^{I}a_{i,j}^{(l)}$. ∎
Appendix B Algorithm for computing the transpose of a partition
The following algorithm computes the transpose of a permutation in exponential notation. Let $a=(a_{l})_{l=1}^{n}$ be the input initial partition and let $(b)=(b_{l})_{l=1}^{n}$ be the transpose of $a$. Suppose that we sort the indices $l$ of $a$ such that $a_{l}\neq 0$. This gives the list $L=(l_{1},l_{2},\ldots,l_{m})$ with $l_{1}>l_{2}>\cdots>l_{m}$ where $m$ is the number of non-zero entries of $a_{l}$. Also set $l_{m+1}=0$. The non-zero entries of $b$ are then
$b_{k_{i}}=l_{i}-l_{i+1}\text{ where }k_{i}=a_{l_{1}}+a_{l_{2}}+\cdots+a_{l_{i}}\text{ for }1\leq i\leq m.$
For example, if $a_{l_{1}}=1$, then the number of $1$’s in $b$ is equal to the difference in size between the two largest parts of $a$. These non-zero values of $b$ can be computed efficiently by going through the list $L$ and updating the partial sums $a_{l_{1}}+a_{l_{2}}\cdots+a_{l_{i}}$. Algorithm 3 implements this idea.
Algorithm 3 Transpose of a partition
1:$(a_{l})_{l=1}^{n}$ (partition of $n$)
2:$(b_{l})_{l=1}^{n}$ (transpose of $(a_{l})_{l=1}^{n}$)
3:$b_{l}\leftarrow 0$ for all $l\in[n]$
4:$L\leftarrow\{l:a_{l}\neq 0\}$
5:Sort $L\leftarrow(l_{1},l_{2},\ldots,l_{m})$ in descending order
6:$k\leftarrow 0$
7:for $i\leftarrow 1$ to $m-1$ do
8:     $k\leftarrow k+a_{l_{i}}$
9:     $b_{k}\leftarrow l_{i}-l_{i+1}$
10:end for
11:$k\leftarrow k+a_{l_{m}}$
12:$b_{k}\leftarrow l_{m}$
13:return $(b_{l})_{l=1}^{n}$
Recall that the number of non-zero elements of $a$ is always at most $2\sqrt{n}$. Thus, the collection $L$ in Algorithm 1 has size at most $2\sqrt{n}$ and the time complexity of sorting $L$ is $O(\sqrt{n}\log n)$. Updating $k$ and computing the difference $l_{i}-l_{i+1}$ has complexity $O(\log n)$. Thus, the complexity of Algorithm 3 is $O(\sqrt{n}\log n)$. Which is comparable to the $O(\sqrt{n}(\log n)^{2})$ complexity of the lumped Burnside process for partitions.